OpenStudy (anonymous):

Stephanie is student teaching at her local high school. She has to help her instructor write a test. She has to write a radical equation where the solution is extraneous. She also has to write a radical equation where the solution is non-extraneous. Help Stephanie write one radical equation where the solution is extraneous and another equation where the solution is non-extraneous. Using complete sentences, explain each step when solving to justify your examples.

OpenStudy (anonymous):

@satellite73 @Loser66 @Compassionate

OpenStudy (anonymous):

@hilbertboy96

OpenStudy (anonymous):

and example of an extraneous radical equation is where one of the answers solved for it can not be included. One example of this is \[x-6 = \sqrt{x}\] because when you solve for x you get \[x^2 - 12x + 36 = x\] \[x^2 - 13x + 36 = 0\] \[(x-9)(x-4) = 0\] \[x = 9, x = 4\] How ever only 9 (meaning it is non-extraneous) works in the original and 4 does not (you can try to verify) Non-extraneous just means all solutions work in the equation. One example could be \[\sqrt{x-4} = 2 \] \[x - 4 = 4\] \[x = 8\] 8 is non-extraneous because if works in the original equation

OpenStudy (anonymous):

As for wording the explanation in complete sentences that is up to you but that is the basic concept.